AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Wed Jun 4 00:10:39 2025
Author
*
:
Subject
*
:
<p>01:10</p> <p>Gas target turned off via the shutter valve.</p> <p>Run 49 started beam only background run.</p> <p>Slow comparator threshold lowered to 0x14 (200 keV) in preparation for the energy change tomorrow morning.</p> <p>Other than that no changes to the DAQ or to the experimental setup.</p> <p>D+ beam remains at 0.5 MeV nominally. All four detectors positioned fully out and stationary,</p> <p>I tried to re-bias detector #4 however was not able to bias past 70V. DSSD4 remains off and set to the MEC (0x1). (see attchement 1)</p> <p>Ring cycle is every 10 seconds. Beam position unchanged, Beam position. Vertical =0.0mm , Horizontal=0.0 mm</p> <p>Ring vacuum is now 3.16e-11 mbar. Before switching off the gas target density was 2e12 at/cm2</p> <p>Data rate is 89 kB/sec</p> <p>ADC data items stats remain largely unchanged at new threshold. (see attachment 2)</p> <p>Aida01 has increased to several thousand. Tried check ASIC control on asic #4 of aida01->no effect on higher rates.</p> <p>01:20</p> <p>All checks OK.</p> <p>After 20 minutes of running see energy histogram and xy histogram (attachments 3,4)</p> <p>Interesting xy pattern (hydrogen? Injection noise?)</p> <p>Note put aida13-aida16 slow comparator threshold back to 0x24 as we had some rates coming through (~1kHz) which is now back to zero. If we revive the detector in the morning, the thresholds should be lowered in line with other detectors.</p> <p>03:00</p> <p>All checks OK</p> <p>04:30</p> <p>All checks ok</p> <p>7:20</p> <p>All checks ok</p> <p><strong>Run stopped</strong>, final energy histogram and xy histogram attached.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
Screenshot_from_2025-06-04_01-18-10.png
Original size: 1920x977
Attachment 2:
Screenshot_from_2025-06-04_01-18-59.png
Original size: 1920x1014
Attachment 3:
Screenshot_from_2025-06-04_01-20-47.png
Original size: 1920x1014
Attachment 4:
Screenshot_from_2025-06-04_01-20-56.png
Original size: 1920x1014
Attachment 5:
Screenshot_from_2025-06-04_07-22-01.png
Original size: 1920x1080
Attachment 6:
Screenshot_from_2025-06-04_07-21-56.png
Original size: 1920x1080
Attachment 7:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.3-7933898